# MobileDiffusion: Subsecond Text-to-Image Generation on Mobile Devices

### v1

@@ 4.3 Applications
Our primary focus is on investigating on-device applications, particularly in the context of lightweight controllable adapters [33, 61] and LoRA fine-tuning. An essential challenge lies in determining whether distilled models or fine-tuned models are better suited for these applications. For both cfg-aware distilled and UFOGen finetuned models, we have explored methods for finetuning them in the realm of controllable text-to-image generation. In the case of translating canny edge to images, we were able to achieve comparable results by finetuning the adapters exclusively. However, for the more complex task of LoRA, we successfully accomplished our goal with the distilled model by fine-tuning only on the distilled timesteps. Figure 4 visually illustrates that MD can generate high-quality samples in the context of controllable text-to-image tasks.
Figure 4: Example applications of MobileDiffusion.
@@ 5 Conclusions
In this paper, we explore pushing the boundary of diffusion model’s efficiency by proposing MobileDiffusion, with the ultimate goal of democratizing text-to-image generation on mobile device. To achieve this goal, we conduct comprehensive studies of the architecture optimization for text-to-image diffusion models, which is a rarely touched area in prior work. Through the studies, we obtained a highly optimized architecture for diffusion UNet, with less than 400 million parameters and more efficient computational operations, while maintaining the quality. Additionally, we apply existing methods for reducing the sampling steps on MobileDiffusion, including progressive distillation and GAN finetuning. With the combined techniques, we are able to achieve the astonishing inference time of less than one second on mobile devices. We also showcase the versatility and utility of MobileDiffusion across a spectrum of text-to-image tasks, featuring its effectiveness for various downstream applications.
Supplementary Material
6 Architecture Optimization
@@
7 Mobile Diffusion Architectures
@@
8 Extended Generated Examples
8.1 Qualitative Examples
In Figure 6, we present an extended comparison between SD-1.5 and our learned MD variants. We have chosen to use SD-1.5 for this comparison since it employs the same text encoder, OpenAI CLIP ViT/L14, as our models. Additionally, we have expanded our selection of prompts to encompass various topics, including style, imagination, and complex compositions, beyond what was originally covered in the main section. This extended comparison underscores the consistent ability of MD to generate visually appealing results and its superiority over SD-1.5 in correctly interpreting specific prompts, such as those related to color and relationship. In Figure 7, we show more photo-realistic generated samples from our 1-step MD model finetuned with UFOGen. In Figure 8, we generate more examples using same prompts as in Figure 7 from our distilled 8-step MD. Beyond applications of canny edge and LoRA control as mentioned in the main context, we also explored more difficult tasks, layout control and layout + LoRA control. We show qualitative results in Figure 9.
Figure 6: Extended qualitative comparison among SD-1.5 and a range of MD variants.
Figure 7: Extended photo-realistic samples from finetuned 1-step MD.
Figure 8: Extended photo-realistic samples from distilled 8-step MD.
Figure 9: Extended applications of MD on layout control and the combination of layout and LoRA style control.

### v2

@@ 5.4 Applications
Our MobileDiffusion framework supports a wide range of downstream conditional generation applications, as illustrated in Figure 4. On tasks including controllable generation, personalized generation and in-painting, our model consistently generate visually appealing results instantly. We want to emphasize that the conditional generation models are trained on the pre-trained diffusion models, and they are seamlessly adapted to the one-step diffusion model without any re-training. Such a flexibility enables our model to be compatible with more potential applications.
Figure 4: MD applications on plugins, LoRA finetuning and inpainting. In each group, top row illustrates the condition or reference images and bottom row shows the results.
@@ 6 Conclusions
In this paper, we explore pushing the boundary of the diffusion model’s efficiency by proposing MobileDiffusion, with the ultimate goal of democratizing text-to-image generation on mobile devices. To achieve this goal, we conduct comprehensive studies of the architecture optimization for text-to-image diffusion models, which is a rarely touched area in prior work. Through the studies, we obtained a highly optimized architecture for diffusion UNet, with less than 400 million parameters and more efficient computational operations, while maintaining the quality. Additionally, we study the adversarial fine-tuning techniques and provide the best recipe for one-step diffusion sampling that excels in both text-to-image generation and various downstream applications. With the combined efforts, we are able to achieve the astonishing inference time of 0.2 seconds on mobile devices.
Appendix 0.A Architecture Optimization
@@
Appendix 0.B Mobile Diffusion Architectures
@@
Appendix 0.C Extended Generated Examples
In this section, we present extended qualitative results of text to image, adapters and LoRA finetuing.
0.C.1 Qualitative Examples
0.C.2 Text to Image
In Figure 6, we select a wide range of prompts to encompass various topics, including style, imagination, and complex compositions, beyond what was originally covered in the main section. This extended comparison underscores the consistent ability of MD to generate visually appealing results in just one-step with a highly efficient architecture.
0.C.3 Adapters and LoRA Finetuing
We can seamlessly adapt pretrained adapters and LoRA weights to one-step diffusion models. In Figure 8 and Figure 7, we compare qualitative results from different settings. It’s noteworthy that original UFOGen diffusion loss outperforms the other two in terms of condition and style faithfulness.
Figure 6: Extended one-step samples from MD.
Figure 7: Ablation on different adversarial finetuning settings on LoRA applications.
Figure 8: Ablation on different adversarial finetuning settings on adapter applications.